OpenStudy (anonymous):

How do you find the y-intercept for point slope (y=mx+b)? Points given: (2000,1500) (2008,2250)

OpenStudy (ajprincess):

y-intercept=b First find m \[m=\frac{y_2-y_1}{x_2-x_1}\] \[m=\frac{2250-1500}{2008-2000}\] Plug in the value of m in y=mx+b to find b. Here y=1500 and x=2000

OpenStudy (anonymous):

@ashdash does that make sense?

OpenStudy (anonymous):

Ok, I did the slope as shown and got 375 over 4 (375/4). Now i plug in 1500=375/4x + B and isolate/solve?

OpenStudy (ajprincess):

put 2000 in place of x in 1500=375/4x + B and then solve for b.

OpenStudy (anonymous):

Okay, I understand it now! Thank you so much! :)

OpenStudy (ajprincess):

u r welcome:)

OpenStudy (anonymous):

@ashdash ... @ajprincess is good help... like medal-deserving help, don't you think ;) ? Glad to see you learn this.... good luck :)

OpenStudy (anonymous):

Yes she really is! I really appreciate the help! :)

OpenStudy (ajprincess):

I am really glad to hear that u hav understood it. That's what is the best reward to me than the medals.:)

OpenStudy (anonymous):

AMEN!

OpenStudy (anonymous):

Medals are nice, but helpers here really just want to see learning... @ashdash, you are doing it right :)

OpenStudy (anonymous):

In that case, I want to give you that best reward @ajprincess ! I just started this account last night and am trying to figure it out as I go. How do i give you a medal? ;)

OpenStudy (anonymous):

heh... since I brought it up, allow me...

OpenStudy (anonymous):

click the blue button labelled "Best Response" beside any one of @ajprincess's responses... that gives her the medal. It's not like it's real gold or anything (unfortunately)... It's not worth anything other than a visible representation that the person is a good helper, but that helps other students see who seems to be respected for their help.

OpenStudy (anonymous):

Oh, okay! Ya'll are awesome! thanks again! :)

OpenStudy (ajprincess):

welcome again:)